OpenStudy (anonymous):

Write the equation of the line in slope-intercept form that passes through the points (-2, 6) and (-1, 4).

OpenStudy (anonymous):

y = -2x + 5 y = 2x - 3 y = 2x + 6 y = -2x + 2